# MAP: Unleashing Hybrid Mamba-Transformer Vision Backbone’s Potential with Masked Autoregressive Pretraining

### v1

@@ 4 Masked Autoregressive Pretraining for Hybrid Backbones
DGCNN (Wang et al., 2019) 31.6 $\pm$ 2.8 40.8 $\pm$ 4.6 19.9 $\pm$ 2.1 16.9 $\pm$ 1.5
Transformer (Vaswani et al., 2017) 87.8 $\pm$ 5.2 93.3 $\pm$ 4.3 84.6 $\pm$ 5.5 89.4 $\pm$ 6.3
Mamba3D (Han et al., 2024) 92.6 $\pm$ 3.7 96.9 $\pm$ 2.4 88.1 $\pm$ 5.3 93.1 $\pm$ 3.6
HybridMT3D 92.8 $\pm$ 3.2 97.0 $\pm$ 1.8 88.4 $\pm$ 4.3 93.1 $\pm$ 3.8
@@
DGCNN+OcCo(Wang et al., 2021) 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
OcCo (Wang et al., 2021) 94.0 $\pm$ 3.6 95.9 $\pm$ 2.7 89.4 $\pm$ 5.1 92.4 $\pm$4.6
PointMamba (Liang et al., 2024) 95.0 $\pm$ 2.3 97.3 $\pm$ 1.8 91.4 $\pm$ 4.4 92.8 $\pm$ 4.0
MaskPoint (Liu et al., 2022a) 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
Point-BERT (Yu et al., 2022) 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
Point-MAE (Pang et al., 2022) 96.3 $\pm$ 2.5 97.8$\pm$ 1.8 92.6 $\pm$4.1 95.0 $\pm$ 3.0
Mamba3d+P-B (Yu et al., 2022) 95.8 $\pm$ 2.7 97.9 $\pm$ 1.4 91.3 $\pm$ 4.7 94.5 $\pm$ 3.3
Mamba3d+P-M (Pang et al., 2022) 96.4 $\pm$ 2.2 98.2 $\pm$1.2 92.4 $\pm$ 4.1 95.2 $\pm$ 2.9
@@
HybridMT3D+MAP 97.3 $\pm$2.8 98.7 $\pm$0.8 93.0 $\pm$ 3.6 96.0 $\pm$ 2.7
@@
PointNet (Qi et al., 2017a) 80.4 83.7 3.6 4.9
PointNet++ (Qi et al., 2017b) 81.9 85.1 1.0 4.9
DGCNN (Wang et al., 2019) 82.3 85.2 1.3 12.4
Transformer (Vaswani et al., 2017) 83.4 85.1 27.1 15.5
Mamba3D(Han et al., 2024) 83.7 85.7 23.0 11.8
HybridMT3D 83.5 85.6 25.1 12.9
@@
OcCo (Wang et al., 2021) 83.4 84.7 27.1 -
PointContrast (Xie et al., 2020) - 85.1 37.9 -
CrossPoint (Afham et al., 2022) - 85.5 - -
Point-MAE (Pang et al., 2022) 84.2 86.1 27.1 15.5
PointMamba (Liang et al., 2024) 84.4 86.0 17.4 14.3
Point-BERT (Yu et al., 2022) 84.1 85.6 27.1 10.6
Mamba3d+P-B  (Yu et al., 2022) 84.1 85.7 21.9 9.5
Mamba3d+P-M (Pang et al., 2022) 84.3 85.8 23.0 11.8
@@
HybridMT3D+MAP 84.7 86.3 25.1 12.9
Table 11: (Left) Few-shot classification on ModelNet40 dataset. (Right) Part segmentation on ShapeNetPart dataset. Our results are highlighted in blue.
5 Experiments
5.1 2D Experiments on ImageNet-1K Classification Task
Settings. We pretrained on the training set of the ImageNet-1K(Deng et al., 2009b) dataset and then fine-tuned on its classification task. We report the top-1 validation accuracy of a single 224x224 crop, and in some settings, we also report the results for a 384x384 crop. During the pretraining phase, we applied a random masking strategy with a 50% masking ratio, using only random cropping as the data augmentation strategy. We utilized AdamW as the optimizer and trained for 1600 epochs across all settings. Additionally, we pretrained using the MAP paradigm on pure Mamba and pure Transformer networks, demonstrating that this paradigm is effective for both frameworks. In the fine-tuning phase, we directly fine-tune for 400 epochs and report the results.
Results. Results are shown in Table 8. The results indicate that the hybrid framework achieves a balance between performance and computational overhead. However, simply training the hybrid architecture from scratch does not lead to significant performance improvements compared to pure Mamba and Transformer backbone. Our proposed pretraining method significantly enhances the performance of the hybrid Mamba-Transformer framework. Additionally, we verified that our MAP method also significantly improves the performance of both the pure Mamba framework and the pure Transformer backbone. Furthermore, when comparing models of the base size with other pretraining methods, we observed that contrastive learning pretraining does not yield performance improvements. The original MAE and AR methods also fail to fully exploit the capabilities of the hybrid Mamba-Transformer backbone, with their results significantly lower than our MAP pretraining method. This further demonstrates the effectiveness of our method for the hybrid framework.
Results with Different Hybrid Ratio for Mamba and Transformer. In our experiments, we used a 3:1 hybrid ratio of Mamba to Transformer. We also explored other hybrid ratios, and the results, as shown in Table 9, indicate that there are no significant performance differences among the hybrid models with varying ratios after MAP pretraining. Considering computational efficiency and memory savings, we opted to adopt the 3:1 hybrid ratio as our default configuration.
5.2 3D Experiments on ModelNet40, ScanObjectNN and ShapeNetPart
Settings. We pretrained using the ShapeNet(Chang et al., 2015) dataset, employing random rotation and translation scaling as data augmentation techniques. Each point cloud consists of 1024 points and is divided into 64 patches, with each patch containing 32 points. We also used a hybrid ratio of Mamba to Transformer at 3:1, randomly masking 50% of the patches. Since point clouds are unordered, the concept of rows does not apply here; instead, we randomly generate 32 patches each time and complete the reconstruction process in an autoregressive manner. Similar to Mamba3DHan et al. (2024), we did not adopt any special sorting strategies but ensured that the order of pretraining matches that of the actual Mamba scans. We conducted pretraining on both the hybrid framework and the original Mamba3D to validate their performance advantages in both the pure Mamba framework and the hybrid framework. During pretraining and downstream fine-tuning, we employed the AdamW optimizer with a cosine decay strategy for 300 epochs. For the ModelNet40Wu et al. (2015) fine-tuning experiments, we used translation and scaling as data augmentation, while on ScanObjectNNUy et al. (2019a), we applied random rotation as data augmentation. Additionally, I also performed experiments in few-shot settings and on ShapeNet part(Yi et al., 2016) segmentation.
Results.The experiments demonstrate that our method significantly enhances the performance of both the hybrid framework and the pure Mamba framework on 3D tasks. This suggests that our approach can be easily adapted to other domains and data types, such as LLMs and video data. Notably, in the part segmentation task, the performance of the hybrid framework trained from scratch is inferior to that of the pure Mamba framework. However, after pretraining, the advantages of the hybrid framework are fully realized, significantly surpassing the performance of the pure Mamba framework. This further proves that our method can simultaneously harness the potential of both Mamba and Transformer to achieve better performance.
6 Conclusion
In this paper, we begin with an in-depth analysis of the key factors that contribute to the success of autoregressive pretraining for Mamba. Based on this, We introduce a pretraining strategy specifically designed for the Mamba-Transformer hybrid framework for the first time. This strategy is effective not only for the hybrid backbones but also for pure Mamba and pure Transformer backbones. We have validated the effectiveness of our approach on both 2D and 3D datasets.

### v2

@@ 6 Conclusion
DGCNN [50] 31.6 $\pm$ 2.8 40.8 $\pm$ 4.6 19.9 $\pm$ 2.1 16.9 $\pm$ 1.5
Transformer [45] 87.8 $\pm$ 5.2 93.3 $\pm$ 4.3 84.6 $\pm$ 5.5 89.4 $\pm$ 6.3
Mamba3D [15] 92.6 $\pm$ 3.7 96.9 $\pm$ 2.4 88.1 $\pm$ 5.3 93.1 $\pm$ 3.6
HybridNet 92.8 $\pm$ 3.2 97.0 $\pm$ 1.8 88.4 $\pm$ 4.3 93.1 $\pm$ 3.8
@@
DGCNN+OcCo[47] 90.6 $\pm$ 2.8 92.5 $\pm$ 1.9 82.9 $\pm$ 1.3 86.5 $\pm$ 2.2
OcCo [47] 94.0 $\pm$ 3.6 95.9 $\pm$ 2.7 89.4 $\pm$ 5.1 92.4 $\pm$4.6
PointMamba [22] 95.0 $\pm$ 2.3 97.3 $\pm$ 1.8 91.4 $\pm$ 4.4 92.8 $\pm$ 4.0
MaskPoint [26] 95.0 $\pm$ 3.7 97.2 $\pm$ 1.7 91.4 $\pm$ 4.0 93.4 $\pm$ 3.5
Point-BERT [56] 94.6 $\pm$ 3.1 96.3 $\pm$ 2.7 91.0 $\pm$ 5.4 92.7 $\pm$ 5.1
Point-MAE [34] 96.3 $\pm$ 2.5 97.8$\pm$ 1.8 92.6 $\pm$4.1 95.0 $\pm$ 3.0
Mamba3d+P-B [56] 95.8 $\pm$ 2.7 97.9 $\pm$ 1.4 91.3 $\pm$ 4.7 94.5 $\pm$ 3.3
Mamba3d+P-M [34] 96.4 $\pm$ 2.2 98.2 $\pm$1.2 92.4 $\pm$ 4.1 95.2 $\pm$ 2.9
@@
HybridNet+MAP 97.3 $\pm$2.8 98.7 $\pm$0.8 93.0 $\pm$ 3.6 96.0 $\pm$ 2.7
Table 12: Few-shot classification on ModelNet40 dataset. The overall accuracy (%) without voting is reported. P-B and P-M represent Point-BERT and Point-MAE strategy, respectively.
Method PT #P $\downarrow$ #F $\downarrow$ ScanObjectNN
OBJ_BG $\uparrow$ OBJ_ONLY $\uparrow$ PB_T50_RS $\uparrow$
Supervised Learning Only: Dedicated Architectures
PointNet[36] $\times$ 3.5 0.5 73.3 79.2 68.0
PointNet++[37] $\times$ 1.5 1.7 82.3 84.3 77.9
DGCNN[50] $\times$ 1.8 2.4 82.8 86.2 78.1
PointCNN[21] $\times$ 0.6 - 86.1 85.5 78.5
DRNet [39] $\times$ - - - - 80.3
SimpleView[10] $\times$ - - - - 80.5$\pm$0.3
GBNet[40] $\times$ 8.8 - - - 81.0
PRA-Ne[6] $\times$ - 2.3 - - 81.0
MVTN[13] $\times$ 11.2 43.7 92.6 92.3 82.8
PointMLP[32] $\times$ 12.6 31.4 - - 85.4$\pm$0.3
PointNeXt[38] $\times$ 1.4 3.6 - - 87.7$\pm$0.4
P2P-HorNet[51] $\checkmark$ - 34.6 - - 89.3
DeLA[4] $\times$ 5.3 1.5 - - 90.4
Supervised Learning Only: Transformer or Mamba-based Models
Transformer $\times$ 22.1 4.8 79.86 80.55 77.24
PCT[12] $\times$ 2.9 2.3 - - -
PointMamba[23] $\times$ 12.3 3.6 88.30 87.78 82.48
PCM[60] $\times$ 34.2 45.0 - - 88.10$\pm$0.3
SPoTr[35] $\times$ 1.7 10.8 - - 88.60
PointConT[28] $\times$ - - - - 90.30
Mamba3d w/o vot.[15] $\times$ 16.9 3.9 92.94 92.08 91.81
Mamba3d w/ vot.[15] $\times$ 16.9 3.9 94.49 92.43 92.64
HybridNet w/o vot. $\times$ 19.3 4.4 92.81 92.28 91.97
HybridNet w/ vot. $\times$ 19.3 4.4 94.50 92.58 92.66
With Self-supervised pretraining
Transformer OcCo 22.1 4.8 84.85 85.54 78.79
Point-BERT IDPT 22.1+1.7 4.8 88.12 88.30 83.69
MaskPoint MaskPoint 22.1 4.8 89.30 88.10 84.30
PointMamba Point-MAE 12.3 3.6 90.71 88.47 84.87
Point-MAE IDPT 22.1+1.7 4.8 91.22 90.02 84.94
Point-M2AE Point-M2AE 15.3 3.6 91.22 88.81 86.43
Mamba3d w/o vot. Point-BERT 16.9 3.9 92.25 91.05 90.11
Point-MAE Point-MAE 22.1 4.8 90.02 88.29 85.18
Mamba3d w/o vot. Point-MAE 16.9 3.9 93.12 92.08 92.05
Mamba3d w/ vot. Point-MAE 16.9 3.9 95.18 94.15 93.05
Mamba3d w/o vot. MAP 16.9 3.9 93.62 92.75 92.65
Mamba3d w/ vot. MAP 16.9 3.9 95.64 94.87 93.76
HybridNet w/o vot. MAP 19.3 4.4 93.88 93.03 92.95
HybridNet w/ vot. MAP 19.3 4.4 95.84 94.97 93.87
Table 13: Results on 3D classification tasks. Our results are highlighted in blue. PT: pre-training strategy.
@@
PointNet [36] 80.4 83.7 3.6 4.9
PointNet++ [37] 81.9 85.1 1.0 4.9
DGCNN [50] 82.3 85.2 1.3 12.4
Transformer [45] 83.4 85.1 27.1 15.5
Mamba3D[15] 83.7 85.7 23.0 11.8
HybridNet 83.5 85.6 25.1 12.9
@@
OcCo [47] 83.4 84.7 27.1 -
PointContrast [54] - 85.1 37.9 -
CrossPoint [1] - 85.5 - -
Point-MAE [34] 84.2 86.1 27.1 15.5
PointMamba [22] 84.4 86.0 17.4 14.3
Point-BERT [56] 84.1 85.6 27.1 10.6
Mamba3d+P-B  [56] 84.1 85.7 21.9 9.5
Mamba3d+P-M [34] 84.3 85.8 23.0 11.8
@@
HybridNet+MAP 84.7 86.3 25.1 12.9
Table 14: Part segmentation on ShapeNetPart dataset. Our results are highlighted in blue. The class mIoU (mIoUC) and the instance mIoU (mIoUI) are reported, with model parameters #P (M) and FLOPs #F (G).